OpenStudy (anonymous):

Which property could be used to justify the transformation from line 2 to line 3 line 1 - 7n= - 2n + 65 line 2 -5n = 65 line 3 N=-13 Distributive property of equality Addition property of equality Multiplication property of equality Symmetric property of equality

OpenStudy (anonymous):

addition

OpenStudy (anonymous):

how are you sure \

OpenStudy (anonymous):

its what i would put

OpenStudy (anonymous):

what about 1 and 2

Directrix (directrix):

Addition is not correct. line 2 -5n = 65 Divide both sides by -5 ==> @joynecia___ line 3 N=-13 That would be Division Property of Equality

Directrix (directrix):

>>what about 1 and 2 line 1) - 7n= - 2n + 65 line 2) -5n = 65 What operation do you apply to -7n to get -5n that you also do to -2n + 65 to get 65?